[Hallc_running] [New Logentry] RC Report -- 09/03/22

dhamilto at jlab.org dhamilto at jlab.org
Sat Sep 3 10:05:02 EDT 2022

Logentry Text:
We are currently in production running at 2.5 uA on LD2 for the final kinematic setting (Q2=1.60, W=3.08 Th_shms=6.00).

Thanks to Dave Mack and our Hall D colleagues MCC were able to take some time yesterday to set up the BCM calibration by doing some studies with the Faraday cup in (https://logbooks.jlab.org/entry/4037104). We will do the second part of the study next week. Otherwise, we have had smooth running since moving to this kinematic setting. CH C1 breaker 2 tripped yesterday around the Day to Swing changeover and Jack reset it (https://logbooks.jlab.org/entry/4037172). We should look at the load on that breaker. The beam was off for several hours during Swing due to problems with a pump in the 2L03 zone. Even so, we got around 1000 e-pi events in the last 24 hours.

Since there is not much to discuss I propose we skip the RC meeting today.


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/4037359
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20220903/14bbc6e1/attachment.html>

More information about the Hallc_running mailing list